# Einstein’s elevator and the principle of equivalence

### v1

@@ Abstract
We outlines here the design, execution, and educational outcomes of an intervention inspired by Einstein’s elevator thought experiment, intended to introduce secondary school students to the principle of equivalence, which is at the basis of the theory of General Relativity. We build an experimental version of Einstein’s elevator, which simulated the effects of free-fall in an accelerated reference frame: a detailed description of the experimental apparatus and its construction is provided, highlighting the challenges and innovations in creating a simple yet functional setup using everyday materials.
@@ I Introduction
The paper is organized as follows: in Section II we describe the development of the intervention, while in Section III we focus on the structure of the lectures and their methodologies. Einstein’s elevator is discussed in Section IV, both in terms of its didactic relevance and with regards to its experimental construction. The results are discussed in Section VI and the conclusions can be found in Section VII.
@@ III.1 Initial test
Before starting the first activity, students were asked to read the already cited excerpt (see Appendix B) taken by the book written by Einstein and Infeld (1966) and answer five multiple-choice questions, which can be found in Appendix A. The last question also required a brief explanation of the chosen answer. The responses collected allowed for an assessment of the student’s level of knowledge before beginning the learning process.
@@
The first four questions in the initial test were designed to encourage students to reflect on the concept of inertial reference frames and the force of gravity, two key themes in Einstein’s theory, which are completely redefined with respect to their meaning in classical physics. The last question, more challenging, asked students to align themselves with Einstein, classical physics, both, or neither, providing justification for their choice. The primary purpose of this question was not so much to elicit the “correct” answer but to encourage students to present a well-reasoned and articulated explanation. An analysis of the answers given by the students will be presented in Section VI.1.
@@ III.4 Formulation of the equivalence principle
During the experiment it was observed that at the moment of free fall, the reading of the dynamometer is zero: this explains why the object or person inside the elevator experiences the sensation of “floating,” as if in a state of weightlessness. As a matter of fact, the weight recorded by the dynamometer is zero. After disproving the first two hypotheses formulated by the students, a collective discussion was initiated to analyze the observations from a theoretical perspective. The sensation of weightlessness and floating inside the elevator, along with the fact that the spring dynamometer registers zero mass, does not guarantee being sufficiently far from a massive body or free from gravitational influences. Building on this reasoning, the principle of equivalence was introduced, encouraging students to articulate it spontaneously rather than presenting it as a “predefined and ready-made law.” To conclude the first two hours of the lesson in an engaging way and to set the stage for the next session, ensuring better assimilation of the concepts, a video was shown featuring the Italian astronaut Samantha Cristoforetti aboard the International Space Station, where she explains how mass and gravity work in space Cristoforetti .
@@ V Projecting the elevator
To simulate the elevator, we decided to use a wooden box and cushion its landing with foam placed on the floor near the door. Inside the box, to represent a person, we would place a 200 g weight attached to a dynamometer, which would be fixed to the ceiling of the box. This setup would allow us to measure fluctuations in weight during the fall. To record the experiment, we planned to design a structure (still undefined at the project’s inception) to hold a smartphone. The box would be suspended by a system of ropes and pulleys that would lift it to the top of the door before releasing it to fall freely to the ground. This arrangement would enable students to conduct the experiment independently and safely.
V.1 Building the elevator
This section outlines the steps involved in building the experimental setup, accompanied by images to enhance clarity. Additionally, extra photographs are in Appendix C to better illustrate the completed setup.
@@ V.1 Building the elevator
As we said, to simulate a person inside the elevator, a weight of 200 g was suspended from a dynamometer with a measurement range of up to 1 kg. The dynamometer was attached to a hook on the upper wall of the box, ensuring that the weight remained suspended without contacting the lower surface FIG. 5. At the base of the door, a wooden plank measuring 75 cm x 12 cm x 1.5 cm was installed FIG. 6. Two hooks were affixed to the plank at intervals that corresponded to the width of the box FIG. 7. This setup served two purposes: first, to cushion the box’s impact by placing foam material between the clamps, and second, to secure the ropes that act as guides for the elevator. To stabilize the ropes and keep the plank in position, weights were added to both ends of the plank FIG. 8.
@@
V.2 Practice tests and modifications
@@
VI Discussion of the Results
VI.1 The initial test
@@ VI.2 During the lesson
VII Conclusions
@@ D. Both
Appendix B The Dialogue with a Classical Physicist
@@
Appendix C Figures

### v2

# thanks: Corresponding Author

@@ Abstract
We outline here the design, execution, and educational outcomes of an intervention inspired by Einstein’s elevator thought experiment, intended to introduce secondary school students to the principle of equivalence, which is at the basis of the theory of General Relativity. We build an experimental version of Einstein’s elevator, which simulated the effects of free-fall in an accelerated reference frame: a detailed description of the experimental apparatus and its construction is provided, highlighting the challenges and innovations in creating a simple yet functional setup using everyday materials.
@@ I Introduction
The paper is organized as follows: in Section II we describe the development of the intervention, while in Section III we focus on the structure of the lectures and their methodologies. Einstein’s elevator is discussed in Section IV, both in terms of its didactic relevance and with regard to its experimental construction. The results are discussed in Section V and the conclusions can be found in Section VI.
@@ III.1 Initial test
Before starting the first activity, students were asked to read the already cited excerpt (see Appendix C.1) taken by the book written by Einstein and Infeld (1966) and answer five multiple-choice questions, which can be found in Appendix A. The last question also required a brief explanation of the chosen answer. The responses collected allowed for an assessment of the student’s level of knowledge before beginning the learning process.
@@
The first four questions in the initial test were designed to encourage students to reflect on the concept of inertial reference frames and the force of gravity, two key themes in Einstein’s theory, which are completely redefined with respect to their meaning in classical physics. The last question, more challenging, asked students to align themselves with Einstein, classical physics, both, or neither, providing justification for their choice. The primary purpose of this question was not so much to elicit the “correct” answer but to encourage students to present a well-reasoned and articulated explanation. An analysis of the answers given by the students will be presented in Section V.1.
@@ III.4 Formulation of the equivalence principle
During the experiment it was observed that at the moment of free fall, the reading of the dynamometer is zero: this explains why the object or person inside the elevator experiences the sensation of “floating,” as if in a state of weightlessness. As a matter of fact, the weight recorded by the dynamometer is zero. After disproving the first two hypotheses formulated by the students, a collective discussion was initiated to analyze the observations from a theoretical perspective. The sensation of weightlessness and floating inside the elevator, along with the fact that the spring dynamometer registers zero weight, does not guarantee being sufficiently far from a massive body or free from gravitational influences. Building on this reasoning, the principle of equivalence was introduced, encouraging students to articulate it spontaneously rather than presenting it as a “predefined and ready-made law.” To conclude the first two hours of the lesson in an engaging way and to set the stage for the next session, ensuring better assimilation of the concepts, a video was shown featuring the Italian astronaut Samantha Cristoforetti aboard the International Space Station, where she explains how mass and gravity work in space Cristoforetti .
@@ IV.1 Projecting the elevator
To simulate the elevator, we decided to use a wooden box and cushion its landing with foam placed on the floor near the door. Inside the box, to represent a person, we would place a 200 g mass attached to a dynamometer, which would be fixed to the ceiling of the box. This setup would allow us to measure fluctuations in weight during the fall. To record the experiment, we planned to design a structure (still undefined at the project’s inception) to hold a smartphone. The box would be suspended by a system of ropes and pulleys that would lift it to the top of the door before releasing it to fall freely to the ground. This arrangement would enable students to conduct the experiment independently and safely.
IV.2 Building the elevator
This section outlines the steps involved in building the experimental setup, accompanied by images to enhance clarity. Additional photographs can be found in Appendix D to better illustrate the completed setup, and a video demonstrating both the apparatus and its operation is available here https://tinyurl.com/elevatorped053025 (2025).
@@ IV.2 Building the elevator
As we said, to simulate a person inside the elevator, a mass of 200 g was suspended from a dynamometer with a measurement range of up to 1 kg. The dynamometer was attached to a hook on the upper wall of the box, ensuring that the weight remained suspended without contacting the lower surface FIG. 5. At the base of the door, a wooden plank measuring 75 cm x 12 cm x 1.5 cm was installed FIG. 6. Two hooks were affixed to the plank at intervals that corresponded to the width of the box FIG. 7. This setup served two purposes: first, to cushion the box’s impact by placing foam material between the clamps, and second, to secure the ropes that act as guides for the elevator. To stabilize the ropes and keep the plank in position, weights were added to both ends of the plank FIG. 8.
@@
IV.3 Practice tests and modifications
@@
V Discussion of the Results
V.1 The initial test
@@ V.2 During the lesson
V.3 Final test
The file containing the questions for the final test (Appendix B) was emailed to the teachers, who subsequently shared it with their students. The test duration was approximately one hour per class, although some classes completed it in as little as 45 minutes. It was administered in paper format, and no personal data from the students was collected. Out of a total of 115 students, 109 participated and answered the final questions.
Due to school scheduling, the final test was administered by the teachers during regular class hours. As a result, a few students who had not attended the lessons also took the test. Although the exact number of these students was not recorded, it is known that it does not exceed five. Their responses were included in the final sample; however, they do not significantly impact the overall results.
Regarding the analysis of the students’ responses to the final test, most of them provided the correct answers to the first four multiple-choice questions. A detailed analysis of the results shows that 90% of the students answered the frst question correctly, 72% answered the second, 73% answered the third, and 77% answered the fourth.
In contrast, the open-ended questions proved to be more challenging and elicited more doubts and uncertainties. This type of question is inherently more complex and requires students to engage in deeper reasoning and demonstrate a greater understanding of specific concepts. The three texts referenced for the open-ended questions were: a document about the Gaia mission from the Italian Space Agency’s website ASI (2013) (Appendix C.2), a text discussing the photon sphere of black holes Wikipedia (2022) (Appendix C.3) and an explanation of how GPS works Wikipedia (2025) (Appendix C.4).
To analyze the results, we categorized the responses to the three open-ended questions as follows:
  1. (A)
    Responses in which students referenced Einstein’s theory and provided reasoning based on the new concepts presented during the four hours of lessons.
  2. (B)
    Responses in which students remained aligned with the theories and concepts of classical physics, offering arguments based on that framework.
  3. (C)
    Responses in which students did not reference any specific theory and provided incomplete answers with weak or often invalid reasoning.
Our primary focus is to determine how many students transitioned from classical physics to Einstein’s theory, indicating their understanding of the latter’s propositions. The data reveals that 44% of the students fell into Category (A), 22% into Category (B), and 34% into Category (C).
VI Conclusions
@@
Ethical Statement: Data used in this research were collected during an undergraduate thesis project. Pupils involved were informed about the treatment of data for research purposes.
@@ D. Both
Appendix B Final Test
  • 1. According to Newton’s proposed theory, gravity is an attractive force present between two bodies with mass. Einstein opposed this definition of gravity; according to him, gravity:
    A. Is not a force.
    B. Is an attractive force but is exerted only by the body with greater mass on the body with lesser mass.
    C. Is an attractive force but is exerted only by the body with lesser mass on the body with greater mass.
    D. Is a force that arises from previous contact between the two bodies involved.
  • 2. Through the free-fall elevator experiment, we observed how an object inside it floats at the moment the elevator falls. This happens because:
    A. We are in an inertial reference frame, and therefore gravity is not present. Since the gravitational force is no longer acting, the object’s weight becomes 0 kg.
    B. The elevator’s acceleration, which is g, causes us to experience an apparent force equal and opposite to the acceleration with which we are attracted to the Earth. These two forces cancel out, making the object float.
    C. The elevator accelerates, but the object inside does not; it remains "stationary" within the elevator, which is why it floats. In other words, the elevator moves relative to the object, but the object does not move.
    D. It is an issue with the experimental apparatus. The distance traveled by the free-falling elevator is very short, and the dynamometer inside oscillates. The fact that the recorded weight is zero is just a coincidence.
  • 3. We have stated Einstein’s Equivalence Principle together. It is related to the fact that:
    A. The Earth orbits the Sun.
    B. Staying still on Earth is equivalent to falling freely inside an elevator.
    C. Gravitational mass and inertial mass are equivalent.
    D. Floating in space and standing still on Earth are two equivalent situations.
  • 4. If a marble is launched horizontally inside a free-falling elevator, it will move only in the horizontal direction. How would you explain this phenomenon?
    A. We are in an inertial reference frame, so gravity does not act, canceling the marble’s motion in the vertical direction.
    B. The motion of the marble is linked to the position of a guide placed horizontally inside the elevator, which makes it move horizontally.
    C. The marble, having mass, also experiences the acceleration g as it moves in free fall. As a result, an apparent force of double intensity and opposite to the force pulling us downward is created, canceling the marble’s motion in the vertical direction.
    D. The marble, having mass, also experiences the acceleration g as it moves in free fall. As a result, an apparent force equal and opposite to the force pulling us downward is created, canceling the marble’s motion in the vertical direction.
    Here are three questions related to the texts: "The Gaia Mission," "The Photon Sphere of Black Holes," and "The GPS." Answer them using the information from the texts and the concepts discussed in class.
  • 5. As explained in the first text, the Gaia mission aims to create a three-dimensional map of our galaxy. Referring to what we discussed in class, explain what potential challenges need to be considered and why General Relativity is essential for this mission.
  • 6. Why is it stated in the second text that light bends near a black hole? Explain in what sense we can say that light curves according to Newton and why it curves according to Einstein.
  • 7. How does the clock on a satellite remain synchronized with the clock on the device being used with GPS?
Appendix C Texts
C.1 The Dialogue with a Classical Physicist
@@
C.2 The Gaia mission
Gaia is a mission of the ESA scientific program aimed at creating a three-dimensional map of our galaxy, revealing its composition, formation, and evolution.
The launch took place on December 19, 2013, using the Soyuz-Fregat launcher, and the satellite was placed in orbit around L2.
The mission will continuously scan the entire sky by leveraging the satellite’s rotation and precession movements: each area of the sky will be observed approximately seventy times during the satellite’s operational lifetime.
The European scientific community’s participation in the mission includes the responsibility of processing the vast amount of data that will be generated. This task is carried out by the Data Processing and Analysis Consortium (DPAC), a consortium of European research institutes established by European scientists in response to an ESA Announcement of Opportunity.
C.3 The photon sphere of the black holes
A photon sphere (or photonic sphere) is a spherical region of space where gravity is strong enough to force photons to move in orbital paths. The formula to determine the radius for a circular photon orbit is: $r=\frac{3GM}{c^{2}}$.
Due to this equation, photon spheres can only exist in the space surrounding an extremely compact object, such as a black hole.
Since photons travel near the event horizon of a black hole, they can escape its gravitational pull by moving in an almost vertical direction, known as the escape cone. A photon at the boundary of this cone will not fully escape the black hole’s gravity; instead, it will orbit around it, although in unstable orbits.
C.4 The GPS
The Global Positioning System (GPS) is a U.S. military satellite-based positioning and navigation system.
Using a dedicated network of artificial satellites in orbit, it provides a mobile terminal or GPS receiver with information about its geographic coordinates and time under any weather conditions, anywhere on Earth or in its immediate vicinity, as long as there is an unobstructed connection with at least four satellites of the system. Localization occurs through the transmission of a radio signal from each satellite and the processing of the received signals by the receiver.
It is managed by the U.S. government and is freely accessible to anyone with a GPS receiver. Its current level of accuracy is within a few meters, depending on weather conditions, the availability and position of the satellites relative to the receiver, the quality and type of receiver, radio signal propagation effects in the ionosphere and troposphere (e.g., refraction), and relativistic effects.
The clocks onboard the satellites are corrected for the effects predicted by relativity, which causes a time shift on the satellites. Observing this shift is considered a real-world confirmation of Einstein’s theory. The detected relativistic effect matches the theoretical expectation within measurement accuracy limits. The shift is the combined result of two factors: due to the satellite’s relative motion, its clock runs 7 microseconds per day slower than clocks on Earth, while due to the weaker gravitational potential at the satellite’s orbit, its clock runs 45 microseconds per day faster. The net result is that the satellite’s clock runs 38 microseconds per day ahead of clocks on Earth.
To compensate for the time difference between satellite and Earth-based clocks, satellite clocks are electronically adjusted. Without these corrections, the GPS system would generate positioning errors of several kilometers per day and would not achieve the centimeter-level accuracy it is capable of.
To reach the indicated accuracy, other synchronization errors must be considered, not just those of relativistic origin. For example, errors related to atmospheric signal propagation or onboard electronics delays can also affect accuracy. While relativistic errors are compensated, correcting atmospheric and electronic errors is more complex.
Appendix D Figures